# Towards Trustworthy Reranking: A Simple yet Effective Abstention Mechanism

### v2

@@ Appendix B Additional Results on Abstention Performance
Dataset normal-$\rightarrow$ SciDocs AskUbuntu StackOverflow Alloprof CMedQAv1 Mmarco
Method normal-$\rightarrow$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$
Model normal-$\downarrow$ Metric normal-$\downarrow$
all-MiniLM-L6-v2 mAP -6.3 73.4 52.5 66.5 6.0 21.6 23.6 13.5 35.6 43.5 19.0 25.1 28.3 64.5 37.1 -8.3 30.5 36.5 28.6 25.0 27.6 36.0 23.2 11.2
mNDCG -2.2 74.9 59.7 71.8 6.7 24.3 29.4 13.5 35.5 43.4 19.3 25.7 27.2 63.1 37.3 -7.0 29.6 36.6 28.6 24.1 27.6 36.1 24.5 11.3
mRR 24.0 91.3 89.8 93.3 13.9 32.5 38.8 21.6 36.5 43.6 20.3 25.6 28.4 61.9 37.9 -7.4 27.9 36.0 28.8 24.1 29.6 37.8 23.3 10.3
all-distilroberta-v1 mAP -5.8 71.3 49.9 63.3 25.2 10.5 26.1 9.0 41.7 47.1 16.5 24.4 17.2 25.3 20.1 17.6 15.5 20.0 3.5 12.4 25.9 23.5 19.7 18.0
mNDCG -0.5 72.0 56.5 67.7 26.5 16.3 31.9 8.3 41.6 47.3 16.7 25.0 16.8 25.0 19.9 17.2 15.4 18.8 4.9 12.4 24.9 21.0 19.8 17.3
mRR 36.8 91.1 93.7 89.1 32.9 24.1 42.2 12.2 42.1 47.4 17.1 24.7 17.0 24.1 19.2 16.9 15.5 18.5 7.9 13.0 26.7 25.2 19.0 18.3
all-mpnet-base-v2 mAP -11.4 74.4 51.0 64.9 19.2 30.3 34.6 19.7 43.4 50.6 20.0 24.1 28.4 40.4 27.2 22.2 19.5 27.8 19.3 22.2 18.9 4.7 7.6 18.8
mNDCG -9.8 75.7 57.0 69.7 22.5 33.7 39.1 17.9 43.4 50.4 20.7 24.8 27.7 40.0 27.6 22.2 18.5 25.7 18.7 21.4 18.7 6.3 8.6 19.8
mRR 10.9 90.2 89.0 93.6 31.5 37.0 45.1 19.1 44.0 50.6 21.6 24.5 28.3 40.4 27.8 22.0 18.4 24.6 19.1 21.3 20.2 7.7 8.2 19.5
bge-base-en-v1.5 mAP -1.4 65.4 46.4 58.2 0.7 38.1 35.2 14.2 46.4 52.8 23.9 32.3 27.8 36.4 17.1 25.0 24.4 37.7 29.1 34.4 38.2 43.8 40.0 27.7
mNDCG 4.7 66.5 53.3 62.1 0.5 38.8 39.2 10.2 46.1 53.1 24.5 32.8 26.8 35.3 17.2 24.8 23.2 36.6 28.8 34.2 36.2 45.2 40.5 27.0
mRR 40.4 83.1 83.7 79.8 3.8 37.6 41.7 3.1 46.4 52.6 24.8 32.5 26.6 34.9 16.6 24.4 23.8 36.2 28.6 34.1 38.5 43.3 39.6 26.6
bge-reranker-base mAP 2.7 58.7 47.1 53.8 -8.9 19.0 24.2 8.1 26.7 35.2 20.7 24.1 38.2 45.0 25.1 30.4 65.1 78.2 74.9 61.4 65.8 34.2 51.3 38.1
mNDCG 8.5 59.2 51.3 57.0 -6.7 21.1 29.4 7.7 26.6 34.9 20.7 24.4 37.7 44.3 25.4 30.8 67.3 79.7 78.1 64.3 64.4 37.3 53.1 40.5
mRR 36.4 66.9 67.0 68.1 0.4 22.9 37.5 13.0 27.5 35.9 20.8 24.3 38.0 44.0 25.3 30.7 70.9 83.6 82.3 67.5 64.6 35.3 53.5 42.3
bge-reranker-large mAP 12.0 65.5 54.2 59.6 -1.7 17.7 23.7 24.4 24.4 33.6 17.1 24.0 44.0 49.0 25.5 32.6 59.6 71.9 70.9 56.9 85.2 86.2 75.8 75.1
mNDCG 18.0 66.9 59.3 64.0 -1.5 15.5 25.3 23.9 24.2 33.6 17.0 24.1 43.6 48.8 25.9 32.9 61.3 75.8 74.5 59.6 85.3 86.4 76.3 75.5
mRR 49.1 79.9 80.4 81.4 4.9 6.1 26.5 24.8 24.6 33.9 17.3 24.2 44.0 48.8 26.0 32.5 64.6 78.5 81.0 64.6 85.2 86.2 75.8 75.1
e5-large-v2 mAP 3.2 71.9 52.0 63.7 12.9 20.4 17.6 16.2 39.2 47.2 23.3 27.6 20.6 33.1 18.3 23.9 22.7 34.0 26.8 30.2 64.2 60.0 52.6 54.0
mNDCG 8.0 72.9 58.1 68.0 16.5 23.9 22.9 15.7 39.1 47.5 23.6 28.1 19.8 32.6 18.4 23.5 22.6 33.4 26.3 29.7 64.0 60.2 53.0 54.2
mRR 41.6 85.8 87.6 88.2 27.1 24.4 27.1 14.1 40.1 48.1 23.9 28.1 19.8 32.3 18.7 22.9 23.9 32.2 24.7 28.2 64.4 59.4 52.4 53.7
e5-small-v2 mAP 1.1 69.2 49.1 61.8 16.6 26.9 23.1 15.8 38.3 47.5 22.4 27.4 20.5 29.4 12.2 17.9 19.9 37.5 30.8 32.9 50.2 48.8 46.7 47.3
mNDCG 7.3 70.7 56.1 66.9 18.7 29.2 26.8 14.3 38.3 47.5 22.8 28.0 20.0 29.0 12.7 17.9 20.3 37.4 30.5 32.9 49.1 46.3 46.3 46.7
mRR 44.2 87.3 86.6 88.3 26.5 29.4 27.6 10.6 38.8 48.1 23.4 27.7 19.8 28.7 13.5 17.4 22.0 37.4 30.8 32.9 50.9 48.5 46.4 46.9
ember-v1 mAP -4.7 73.6 50.7 65.2 10.2 28.6 30.2 16.3 48.0 54.6 23.7 31.7 20.8 27.6 16.4 20.4 23.6 35.9 28.2 30.5 37.9 41.3 42.2 36.4
mNDCG 0.4 73.8 56.8 68.4 14.2 30.8 35.2 13.9 47.8 54.0 24.3 32.3 20.2 27.3 16.9 20.4 22.3 34.3 27.5 29.7 36.8 40.4 41.4 35.8
mRR 34.6 83.6 83.4 81.1 22.2 31.0 41.1 11.3 48.1 54.0 25.0 32.0 20.8 26.9 17.0 19.8 21.7 33.6 27.2 30.6 36.8 41.5 42.1 36.8
llm-embedder mAP 1.9 69.9 47.2 62.3 1.7 31.6 31.7 20.7 47.1 52.6 26.0 32.3 22.8 34.5 19.7 22.7 21.5 36.7 22.4 29.8 40.5 24.7 18.0 25.5
mNDCG 5.8 69.9 52.9 66.4 2.1 34.7 36.7 19.4 47.0 52.7 26.5 32.8 22.5 34.4 20.0 22.6 20.6 36.0 23.0 29.9 40.6 25.5 18.8 25.4
mRR 36.7 86.4 83.1 85.4 6.4 40.3 42.0 19.8 47.8 53.2 26.8 32.5 23.0 34.5 19.7 22.5 21.6 37.5 25.1 31.4 39.8 25.7 19.0 26.4
ms-marco-MiniLM-L-6-v2 mAP 16.1 65.1 52.6 56.3 5.0 20.1 21.7 24.3 38.3 44.3 18.8 21.3 19.9 34.4 24.7 29.3 13.2 16.4 9.2 4.1 18.7 30.1 -6.8 32.9
mNDCG 25.1 66.6 58.8 61.9 3.6 23.1 25.9 27.3 38.0 44.4 19.2 21.9 19.4 34.1 24.7 29.0 13.4 16.3 9.2 3.5 17.3 30.8 -4.9 32.4
mRR 72.4 86.4 87.4 86.9 6.2 20.6 29.5 32.7 38.6 44.6 19.4 21.9 20.2 33.9 25.0 29.1 14.5 16.6 8.1 4.5 19.7 28.1 -7.3 32.4
ms-marco-TinyBERT-L-2-v2 mAP 27.6 60.7 53.6 55.6 0.6 18.0 26.5 27.3 32.2 41.7 24.4 20.0 14.0 24.0 12.1 7.9 -2.3 8.9 5.4 -9.1 29.0 30.3 26.5 45.0
mNDCG 35.4 62.9 59.2 61.0 -0.3 22.4 29.4 29.0 31.9 41.5 24.5 20.5 13.6 23.8 12.4 7.9 -3.3 8.0 4.7 -8.9 27.4 29.5 26.8 43.5
mRR 76.1 83.4 85.8 85.4 2.4 21.9 30.3 28.4 32.4 41.7 24.5 20.2 14.1 23.8 12.2 8.0 -2.9 5.5 3.4 -7.5 29.8 31.2 27.2 44.2
msmarco-MiniLM-L6-cos-v5 mAP 1.4 66.3 51.5 59.9 10.1 19.3 19.5 12.0 36.0 40.7 22.1 22.1 15.1 20.3 10.6 10.0 16.9 18.2 11.0 9.9 49.3 60.8 21.9 26.8
mNDCG 7.6 67.4 57.0 63.1 10.3 22.4 24.3 10.1 36.0 41.0 22.6 22.7 14.1 19.3 10.8 9.8 16.1 17.6 10.5 9.0 48.2 58.1 21.9 26.3
mRR 40.4 79.7 79.5 74.9 14.9 24.3 28.4 6.6 36.6 41.8 23.4 23.3 14.0 17.8 11.1 9.8 17.0 18.5 10.2 7.8 48.7 59.6 22.1 25.4
msmarco-distilbert-dot-v5 mAP 2.3 66.1 50.3 59.7 2.1 15.2 15.1 13.0 36.4 40.2 20.8 24.3 8.7 22.5 11.4 16.2 6.8 5.5 1.3 8.4 54.6 32.1 -0.1 21.8
mNDCG 8.0 67.4 55.5 63.3 3.0 17.9 20.1 13.1 36.3 40.1 21.1 24.7 8.4 22.2 11.8 15.6 6.4 5.2 2.1 8.2 53.6 29.8 1.0 20.6
mRR 41.0 79.5 77.7 76.4 11.1 24.9 26.6 13.7 37.2 40.9 21.2 24.6 8.8 22.1 11.4 15.9 7.5 6.2 3.6 6.1 53.4 28.8 0.2 19.9
multi-qa-MiniLM-L6-cos-v1 mAP -2.1 68.5 52.7 60.6 13.0 13.7 25.1 6.7 42.0 47.5 19.7 27.1 19.0 35.3 32.4 23.2 19.7 34.9 26.6 30.4 22.1 -3.0 15.3 21.6
mNDCG 4.6 70.5 59.6 66.1 14.6 16.1 30.7 3.0 41.8 47.5 20.2 27.5 18.4 34.9 32.3 22.8 18.5 34.7 27.1 30.4 22.2 -4.4 15.3 21.1
mRR 40.4 86.9 88.2 87.6 19.2 23.4 40.2 1.6 42.5 47.8 20.8 27.1 18.8 35.2 33.0 23.3 18.8 35.4 28.5 31.5 22.2 -0.2 15.5 22.4
multi-qa-distilbert-cos-v1 mAP -1.4 68.1 52.3 60.5 20.7 31.9 32.9 27.7 38.6 45.9 19.4 24.1 21.2 34.6 27.0 24.6 20.4 25.3 12.1 18.3 40.5 20.5 29.0 30.9
mNDCG 4.3 68.8 58.6 65.2 22.6 32.5 38.0 25.8 38.6 46.0 19.6 24.7 20.7 34.4 27.5 24.7 19.8 25.8 12.7 17.9 39.4 21.8 28.8 30.7
mRR 35.9 81.9 83.7 82.5 27.1 35.0 46.5 26.0 39.4 46.3 20.1 24.5 20.4 34.0 27.6 25.1 21.5 27.8 14.6 18.0 40.6 20.0 29.0 30.5
multilingual-e5-large mAP 2.8 69.5 49.4 62.6 -16.2 16.8 17.2 10.5 43.2 50.2 23.0 27.5 37.5 43.1 16.9 27.7 48.9 62.0 59.3 47.3 87.2 69.6 74.3 52.7
mNDCG 7.8 70.9 55.8 66.9 -16.3 18.7 21.5 10.7 43.0 50.3 23.4 28.0 37.1 42.7 17.4 27.8 51.1 65.3 63.5 49.4 86.8 68.8 75.2 54.3
mRR 40.7 86.2 88.0 86.7 -13.8 17.0 24.9 13.2 43.8 50.8 23.7 27.8 37.1 42.7 18.1 27.1 56.7 70.0 70.2 52.0 87.3 69.4 76.5 57.1
multilingual-e5-small mAP 4.1 67.6 46.1 59.3 -0.6 28.0 29.8 20.3 45.5 52.9 25.4 31.4 20.4 29.5 8.5 22.3 48.3 68.1 61.1 56.1 86.0 53.0 62.8 43.5
mNDCG 9.1 68.6 52.2 63.7 1.2 33.7 34.1 22.9 45.4 53.0 25.7 31.8 19.8 28.7 8.7 22.2 50.1 70.1 63.7 57.5 86.0 52.8 63.5 44.6
mRR 41.8 84.9 82.7 86.4 5.5 37.0 37.6 29.7 45.8 53.6 25.5 31.4 19.7 27.2 9.3 21.1 54.0 73.9 67.8 59.5 85.9 53.0 62.9 43.3
qnli-distilroberta-base mAP 10.7 25.8 19.3 23.8 3.9 14.1 11.2 8.5 8.9 17.3 13.8 12.7 -1.1 -3.2 2.2 0.5 -1.7 -2.9 -5.5 2.3 1.2 1.3 17.6 8.3
mNDCG 13.2 25.6 19.0 23.0 4.6 14.7 10.3 6.8 8.1 16.9 13.6 12.5 -1.0 -2.2 2.1 0.8 -1.5 -3.6 -4.8 2.0 1.7 2.6 18.1 7.9
mRR 18.8 24.9 18.6 21.2 2.8 15.1 8.9 6.8 8.0 16.0 13.6 12.0 -0.7 -1.9 1.5 0.6 0.2 -3.7 -3.4 1.3 2.2 2.3 16.6 8.0
quora-distilroberta-base mAP 56.3 61.5 60.7 61.0 25.8 26.3 27.6 27.6 20.5 21.9 18.0 17.7 -1.9 5.7 -1.3 -0.9 6.3 4.4 7.9 6.8 3.8 4.7 20.2 25.6
mNDCG 60.9 65.0 64.5 64.8 29.2 30.6 31.1 31.2 20.4 21.9 17.9 17.7 -1.4 4.9 -1.6 -1.5 6.0 5.4 8.4 6.8 2.8 5.7 18.9 24.1
mRR 75.8 76.7 76.6 76.5 31.1 32.5 32.6 32.8 20.0 21.2 17.6 17.3 -1.0 3.6 -1.2 -1.2 6.5 6.9 9.8 7.1 4.4 4.6 19.4 24.5
stsb-TinyBERT-L-4 mAP -8.8 61.0 45.2 56.4 7.3 20.0 14.7 13.6 25.7 33.5 16.3 15.9 -1.1 2.3 7.2 -1.8 1.5 6.1 -2.0 3.3 21.0 13.8 13.0 22.5
mNDCG -4.0 62.2 51.2 60.5 9.3 18.6 16.4 12.2 25.3 33.3 16.6 16.3 -1.9 2.5 7.7 -2.3 1.5 6.0 -1.3 4.0 20.0 17.1 12.2 21.1
mRR 20.7 73.5 73.2 73.6 17.6 18.3 21.4 12.9 25.7 33.5 17.7 16.6 -1.8 3.4 8.0 -2.6 0.2 7.9 -1.3 5.4 20.1 17.1 12.0 21.2
stsb-distilroberta-base mAP 3.8 65.7 49.2 56.8 6.5 16.9 19.3 7.6 28.1 36.2 21.8 23.1 4.7 18.7 -4.7 11.2 11.1 9.7 15.0 -0.5 -1.1 31.5 11.4 -0.8
mNDCG 11.4 66.9 55.0 61.8 8.8 20.2 21.1 7.3 27.9 36.6 21.8 23.5 4.5 18.9 -4.9 10.5 10.3 9.6 14.1 -1.8 0.2 32.7 11.2 -0.9
mRR 50.4 80.7 78.6 80.4 14.2 20.7 23.9 12.2 28.6 36.9 22.1 23.5 4.1 18.0 -5.7 9.1 10.8 8.7 15.6 -1.8 -0.4 31.0 8.8 -2.3

### v3

@@ Appendix B Additional Results on Abstention Performance
Dataset $\rightarrow$ SciDocs AskUbuntu StackOverflow Alloprof CMedQAv1 Mmarco
Method $\rightarrow$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$
Model $\downarrow$ Metric $\downarrow$
ember-v1 mAP 50.7 65.2 -4.7 73.6 30.2 16.3 10.2 28.6 23.7 31.7 48.0 54.6 16.4 20.4 20.8 27.6 28.2 30.5 23.6 35.9 42.2 36.4 37.9 41.3
mNDCG 56.8 68.4 0.4 73.8 35.2 13.9 14.2 30.8 24.3 32.3 47.8 54.0 16.9 20.4 20.2 27.3 27.5 29.7 22.3 34.3 41.4 35.8 36.8 40.4
mRR 83.4 81.1 34.6 83.6 41.1 11.3 22.2 31.0 25.0 32.0 48.1 54.0 17.0 19.8 20.8 26.9 27.2 30.6 21.7 33.6 42.1 36.8 36.8 41.5
llm-embedder mAP 47.2 62.3 1.9 69.9 31.7 20.7 1.7 31.6 26.0 32.3 47.1 52.6 19.7 22.7 22.8 34.5 22.4 29.8 21.5 36.7 18.0 25.5 40.5 24.7
mNDCG 52.9 66.4 5.8 69.9 36.7 19.4 2.1 34.7 26.5 32.8 47.0 52.7 20.0 22.6 22.5 34.4 23.0 29.9 20.6 36.0 18.8 25.4 40.6 25.5
mRR 83.1 85.4 36.7 86.4 42.0 19.8 6.4 40.3 26.8 32.5 47.8 53.2 19.7 22.5 23.0 34.5 25.1 31.4 21.6 37.5 19.0 26.4 39.8 25.7
bge-base-en-v1.5 mAP 46.4 58.2 -1.4 65.4 35.2 14.2 0.7 38.1 23.9 32.3 46.4 52.8 17.1 25.0 27.8 36.4 29.1 34.4 24.4 37.7 40.0 27.7 38.2 43.8
mNDCG 53.3 62.1 4.7 66.5 39.2 10.2 0.5 38.8 24.5 32.8 46.1 53.1 17.2 24.8 26.8 35.3 28.8 34.2 23.2 36.6 40.5 27.0 36.2 45.2
mRR 83.7 79.8 40.4 83.1 41.7 3.1 3.8 37.6 24.8 32.5 46.4 52.6 16.6 24.4 26.6 34.9 28.6 34.1 23.8 36.2 39.6 26.6 38.5 43.3
bge-reranker-base mAP 47.1 53.8 2.7 58.7 24.2 8.1 -8.9 19.0 20.7 24.1 26.7 35.2 25.1 30.4 38.2 45.0 74.9 61.4 65.1 78.2 51.3 38.1 65.8 34.2
mNDCG 51.3 57.0 8.5 59.2 29.4 7.7 -6.7 21.1 20.7 24.4 26.6 34.9 25.4 30.8 37.7 44.3 78.1 64.3 67.3 79.7 53.1 40.5 64.4 37.3
mRR 67.0 68.1 36.4 66.9 37.5 13.0 0.4 22.9 20.8 24.3 27.5 35.9 25.3 30.7 38.0 44.0 82.3 67.5 70.9 83.6 53.5 42.3 64.6 35.3
bge-reranker-large mAP 54.2 59.6 12.0 65.5 23.7 24.4 -1.7 17.7 17.1 24.0 24.4 33.6 25.5 32.6 44.0 49.0 70.9 56.9 59.6 71.9 75.8 75.1 85.2 86.2
mNDCG 59.3 64.0 18.0 66.9 25.3 23.9 -1.5 15.5 17.0 24.1 24.2 33.6 25.9 32.9 43.6 48.8 74.5 59.6 61.3 75.8 76.3 75.5 85.3 86.4
mRR 80.4 81.4 49.1 79.9 26.5 24.8 4.9 6.1 17.3 24.2 24.6 33.9 26.0 32.5 44.0 48.8 81.0 64.6 64.6 78.5 75.8 75.1 85.2 86.2
e5-small-v2 mAP 49.1 61.8 1.1 69.2 23.1 15.8 16.6 26.9 22.4 27.4 38.3 47.5 12.2 17.9 20.5 29.4 30.8 32.9 19.9 37.5 46.7 47.3 50.2 48.8
mNDCG 56.1 66.9 7.3 70.7 26.8 14.3 18.7 29.2 22.8 28.0 38.3 47.5 12.7 17.9 20.0 29.0 30.5 32.9 20.3 37.4 46.3 46.7 49.1 46.3
mRR 86.6 88.3 44.2 87.3 27.6 10.6 26.5 29.4 23.4 27.7 38.8 48.1 13.5 17.4 19.8 28.7 30.8 32.9 22.0 37.4 46.4 46.9 50.9 48.5
e5-large-v2 mAP 52.0 63.7 3.2 71.9 17.6 16.2 12.9 20.4 23.3 27.6 39.2 47.2 18.3 23.9 20.6 33.1 26.8 30.2 22.7 34.0 52.6 54.0 64.2 60.0
mNDCG 58.1 68.0 8.0 72.9 22.9 15.7 16.5 23.9 23.6 28.1 39.1 47.5 18.4 23.5 19.8 32.6 26.3 29.7 22.6 33.4 53.0 54.2 64.0 60.2
mRR 87.6 88.2 41.6 85.8 27.1 14.1 27.1 24.4 23.9 28.1 40.1 48.1 18.7 22.9 19.8 32.3 24.7 28.2 23.9 32.2 52.4 53.7 64.4 59.4
multilingual-e5-small mAP 46.1 59.3 4.1 67.6 29.8 20.3 -0.6 28.0 25.4 31.4 45.5 52.9 8.5 22.3 20.4 29.5 61.1 56.1 48.3 68.1 62.8 43.5 86.0 53.0
mNDCG 52.2 63.7 9.1 68.6 34.1 22.9 1.2 33.7 25.7 31.8 45.4 53.0 8.7 22.2 19.8 28.7 63.7 57.5 50.1 70.1 63.5 44.6 86.0 52.8
mRR 82.7 86.4 41.8 84.9 37.6 29.7 5.5 37.0 25.5 31.4 45.8 53.6 9.3 21.1 19.7 27.2 67.8 59.5 54.0 73.9 62.9 43.3 85.9 53.0
multilingual-e5-large mAP 49.4 62.6 2.8 69.5 17.2 10.5 -16.2 16.8 23.0 27.5 43.2 50.2 16.9 27.7 37.5 43.1 59.3 47.3 48.9 62.0 74.3 52.7 87.2 69.6
mNDCG 55.8 66.9 7.8 70.9 21.5 10.7 -16.3 18.7 23.4 28.0 43.0 50.3 17.4 27.8 37.1 42.7 63.5 49.4 51.1 65.3 75.2 54.3 86.8 68.8
mRR 88.0 86.7 40.7 86.2 24.9 13.2 -13.8 17.0 23.7 27.8 43.8 50.8 18.1 27.1 37.1 42.7 70.2 52.0 56.7 70.0 76.5 57.1 87.3 69.4
msmarco-MiniLM-L6-cos-v5 mAP 51.5 59.9 1.4 66.3 19.5 12.0 10.1 19.3 22.1 22.1 36.0 40.7 10.6 10.0 15.1 20.3 11.0 9.9 16.9 18.2 21.9 26.8 49.3 60.8
mNDCG 57.0 63.1 7.6 67.4 24.3 10.1 10.3 22.4 22.6 22.7 36.0 41.0 10.8 9.8 14.1 19.3 10.5 9.0 16.1 17.6 21.9 26.3 48.2 58.1
mRR 79.5 74.9 40.4 79.7 28.4 6.6 14.9 24.3 23.4 23.3 36.6 41.8 11.1 9.8 14.0 17.8 10.2 7.8 17.0 18.5 22.1 25.4 48.7 59.6
msmarco-distilbert-dot-v5 mAP 50.3 59.7 2.3 66.1 15.1 13.0 2.1 15.2 20.8 24.3 36.4 40.2 11.4 16.2 8.7 22.5 1.3 8.4 6.8 5.5 -0.1 21.8 54.6 32.1
mNDCG 55.5 63.3 8.0 67.4 20.1 13.1 3.0 17.9 21.1 24.7 36.3 40.1 11.8 15.6 8.4 22.2 2.1 8.2 6.4 5.2 1.0 20.6 53.6 29.8
mRR 77.7 76.4 41.0 79.5 26.6 13.7 11.1 24.9 21.2 24.6 37.2 40.9 11.4 15.9 8.8 22.1 3.6 6.1 7.5 6.2 0.2 19.9 53.4 28.8
ms-marco-TinyBERT-L-2-v2 mAP 53.6 55.6 27.6 60.7 26.5 27.3 0.6 18.0 24.4 20.0 32.2 41.7 12.1 7.9 14.0 24.0 5.4 -9.1 -2.3 8.9 26.5 45.0 29.0 30.3
mNDCG 59.2 61.0 35.4 62.9 29.4 29.0 -0.3 22.4 24.5 20.5 31.9 41.5 12.4 7.9 13.6 23.8 4.7 -8.9 -3.3 8.0 26.8 43.5 27.4 29.5
mRR 85.8 85.4 76.1 83.4 30.3 28.4 2.4 21.9 24.5 20.2 32.4 41.7 12.2 8.0 14.1 23.8 3.4 -7.5 -2.9 5.5 27.2 44.2 29.8 31.2
ms-marco-MiniLM-L-6-v2 mAP 52.6 56.3 16.1 65.1 21.7 24.3 5.0 20.1 18.8 21.3 38.3 44.3 24.7 29.3 19.9 34.4 9.2 4.1 13.2 16.4 -6.8 32.9 18.7 30.1
mNDCG 58.8 61.9 25.1 66.6 25.9 27.3 3.6 23.1 19.2 21.9 38.0 44.4 24.7 29.0 19.4 34.1 9.2 3.5 13.4 16.3 -4.9 32.4 17.3 30.8
mRR 87.4 86.9 72.4 86.4 29.5 32.7 6.2 20.6 19.4 21.9 38.6 44.6 25.0 29.1 20.2 33.9 8.1 4.5 14.5 16.6 -7.3 32.4 19.7 28.1
stsb-TinyBERT-L-4 mAP 45.2 56.4 -8.8 61.0 14.7 13.6 7.3 20.0 16.3 15.9 25.7 33.5 7.2 -1.8 -1.1 2.3 -2.0 3.3 1.5 6.1 13.0 22.5 21.0 13.8
mNDCG 51.2 60.5 -4.0 62.2 16.4 12.2 9.3 18.6 16.6 16.3 25.3 33.3 7.7 -2.3 -1.9 2.5 -1.3 4.0 1.5 6.0 12.2 21.1 20.0 17.1
mRR 73.2 73.6 20.7 73.5 21.4 12.9 17.6 18.3 17.7 16.6 25.7 33.5 8.0 -2.6 -1.8 3.4 -1.3 5.4 0.2 7.9 12.0 21.2 20.1 17.1
stsb-distilroberta-base mAP 49.2 56.8 3.8 65.7 19.3 7.6 6.5 16.9 21.8 23.1 28.1 36.2 -4.7 11.2 4.7 18.7 15.0 -0.5 11.1 9.7 11.4 -0.8 -1.1 31.5
mNDCG 55.0 61.8 11.4 66.9 21.1 7.3 8.8 20.2 21.8 23.5 27.9 36.6 -4.9 10.5 4.5 18.9 14.1 -1.8 10.3 9.6 11.2 -0.9 0.2 32.7
mRR 78.6 80.4 50.4 80.7 23.9 12.2 14.2 20.7 22.1 23.5 28.6 36.9 -5.7 9.1 4.1 18.0 15.6 -1.8 10.8 8.7 8.8 -2.3 -0.4 31.0
multi-qa-distilbert-cos-v1 mAP 52.3 60.5 -1.4 68.1 32.9 27.7 20.7 31.9 19.4 24.1 38.6 45.9 27.0 24.6 21.2 34.6 12.1 18.3 20.4 25.3 29.0 30.9 40.5 20.5
mNDCG 58.6 65.2 4.3 68.8 38.0 25.8 22.6 32.5 19.6 24.7 38.6 46.0 27.5 24.7 20.7 34.4 12.7 17.9 19.8 25.8 28.8 30.7 39.4 21.8
mRR 83.7 82.5 35.9 81.9 46.5 26.0 27.1 35.0 20.1 24.5 39.4 46.3 27.6 25.1 20.4 34.0 14.6 18.0 21.5 27.8 29.0 30.5 40.6 20.0
multi-qa-MiniLM-L6-cos-v1 mAP 52.7 60.6 -2.1 68.5 25.1 6.7 13.0 13.7 19.7 27.1 42.0 47.5 32.4 23.2 19.0 35.3 26.6 30.4 19.7 34.9 15.3 21.6 22.1 -3.0
mNDCG 59.6 66.1 4.6 70.5 30.7 3.0 14.6 16.1 20.2 27.5 41.8 47.5 32.3 22.8 18.4 34.9 27.1 30.4 18.5 34.7 15.3 21.1 22.2 -4.4
mRR 88.2 87.6 40.4 86.9 40.2 1.6 19.2 23.4 20.8 27.1 42.5 47.8 33.0 23.3 18.8 35.2 28.5 31.5 18.8 35.4 15.5 22.4 22.2 -0.2
all-MiniLM-L6-v2 mAP 52.5 66.5 -6.3 73.4 23.6 13.5 6.0 21.6 19.0 25.1 35.6 43.5 37.1 -8.3 28.3 64.5 28.6 25.0 30.5 36.5 23.2 11.2 27.6 36.0
mNDCG 59.7 71.8 -2.2 74.9 29.4 13.5 6.7 24.3 19.3 25.7 35.5 43.4 37.3 -7.0 27.2 63.1 28.6 24.1 29.6 36.6 24.5 11.3 27.6 36.1
mRR 89.8 93.3 24.0 91.3 38.8 21.6 13.9 32.5 20.3 25.6 36.5 43.6 37.9 -7.4 28.4 61.9 28.8 24.1 27.9 36.0 23.3 10.3 29.6 37.8
all-distilroberta-v1 mAP 49.9 63.3 -5.8 71.3 26.1 9.0 25.2 10.5 16.5 24.4 41.7 47.1 20.1 17.6 17.2 25.3 3.5 12.4 15.5 20.0 19.7 18.0 25.9 23.5
mNDCG 56.5 67.7 -0.5 72.0 31.9 8.3 26.5 16.3 16.7 25.0 41.6 47.3 19.9 17.2 16.8 25.0 4.9 12.4 15.4 18.8 19.8 17.3 24.9 21.0
mRR 93.7 89.1 36.8 91.1 42.2 12.2 32.9 24.1 17.1 24.7 42.1 47.4 19.2 16.9 17.0 24.1 7.9 13.0 15.5 18.5 19.0 18.3 26.7 25.2
all-mpnet-base-v2 mAP 51.0 64.9 -11.4 74.4 34.6 19.7 19.2 30.3 20.0 24.1 43.4 50.6 27.2 22.2 28.4 40.4 19.3 22.2 19.5 27.8 7.6 18.8 18.9 4.7
mNDCG 57.0 69.7 -9.8 75.7 39.1 17.9 22.5 33.7 20.7 24.8 43.4 50.4 27.6 22.2 27.7 40.0 18.7 21.4 18.5 25.7 8.6 19.8 18.7 6.3
mRR 89.0 93.6 10.9 90.2 45.1 19.1 31.5 37.0 21.6 24.5 44.0 50.6 27.8 22.0 28.3 40.4 19.1 21.3 18.4 24.6 8.2 19.5 20.2 7.7
quora-distilroberta-base mAP 60.7 61.0 56.3 61.5 27.6 27.6 25.8 26.3 18.0 17.7 20.5 21.9 -1.3 -0.9 -1.9 5.7 7.9 6.8 6.3 4.4 20.2 25.6 3.8 4.7
mNDCG 64.5 64.8 60.9 65.0 31.1 31.2 29.2 30.6 17.9 17.7 20.4 21.9 -1.6 -1.5 -1.4 4.9 8.4 6.8 6.0 5.4 18.9 24.1 2.8 5.7
mRR 76.6 76.5 75.8 76.7 32.6 32.8 31.1 32.5 17.6 17.3 20.0 21.2 -1.2 -1.2 -1.0 3.6 9.8 7.1 6.5 6.9 19.4 24.5 4.4 4.6
qnli-distilroberta-base mAP 19.3 23.8 10.7 25.8 11.2 8.5 3.9 14.1 13.8 12.7 8.9 17.3 2.2 0.5 -1.1 -3.2 -5.5 2.3 -1.7 -2.9 17.6 8.3 1.2 1.3
mNDCG 19.0 23.0 13.2 25.6 10.3 6.8 4.6 14.7 13.6 12.5 8.1 16.9 2.1 0.8 -1.0 -2.2 -4.8 2.0 -1.5 -3.6 18.1 7.9 1.7 2.6
mRR 18.6 21.2 18.8 24.9 8.9 6.8 2.8 15.1 13.6 12.0 8.0 16.0 1.5 0.6 -0.7 -1.9 -3.4 1.3 0.2 -3.7 16.6 8.0 2.2 2.3